What is the probability of drawing a white ball from an urn containing 3 white and 4 black balls after having been transferred to it two balls drawn at random from another urn containing 4 white and 4 black balls?
1
Expert's answer
2020-07-09T19:45:40-0400

Solution:

Let A1 be event that  "two balls drawn at random from urn containing 4 white and 4 black balls are white".

Then P(A1)=m/n

m=binomial(4,2)=4!/2!/2!=6

n=binomial(8,2)=8!/2!/6!=28

P(A1)=6/28=3/14


Let A2 be event of "drawing one white ball and one black ball from urn containing 4 white and 4 black balls".

Then P(A2)=m/n

m=binomial(4,1)*binomial(4,1)=4*4=16

n=binomial(8,2)=8!/2!/6!=28

P(A2)=16/28=4/7.


Let A3 be event that "two balls drawn at random from urn containing 4 white and 4 black balls are black".

Then P(A3)=m/n

m=binomial(4,2)=4!/2!/2!=6

n=binomial(8,2)=8!/2!/6!=28

P(A3)=6/28=3/14.


"A_1\bigcup A_2 \bigcup A_3 = U,
A_i\bigcap A_j =\varnothing, 1\le i<j\le 3"


Let B be event of "drawing a white ball from an urn containing 3 white and 4 black balls after having been transferred to it two balls drawn at random from another urn containing 4 white and 4 black balls"

Then

P(B)=P(B|A1)P(A1)+P(B|A2)P(A2)+P(B|A3)P(A3).


P(B|A1)=m/n

m=5, n=9

P(B|A1)=5/9


P(B|A2)=m/n

m=4, n=9

P(B|A2)=4/9


P(B|A3)=m/n

m=3, n=9

P(B|A3)=3/9=1/3


P(B)=5/9*3/14+4/9*4/7+1/3*3/14=4/9


Answer: the probability of drawing a white ball from an urn containing 3 white and 4 black balls after having been transferred to it two balls drawn at random from another urn containing 4 white and 4 black balls is equel 4/9